# Grassroots Consensus

### v2

@@ 3.2 Liveness
Next, we discuss liveness. We first prove that Grassroots Consensus disseminates all correct agents’ blocks and the blocks they depend on.
@@
If a block $b$ is issued by a correct agent in Grassroots Consensus then $[b]$ is eventually received by all correct agents.
Proof 3.11.
By way of contradiction assume that $p,q\in P$ are correct, $b_{p}$ is issued by $p$ (and received by $q$ since the network is reliable) and some block $b\in[b_{p}]$ is never received by $q$. Since by assumption $b$ (and possibly other blocks in $[b_{p}]$) have not been received by $q$, then upon receipt of $b_{p}$, $q$ issues a nack-block $b_{q}$ for which $b\in[b_{q}]$. Upon receipt of $b_{q}$ by $p$, $p$ will send (once) $[b_{q}]$, which includes $b$, to $q$. Since the network is reliable, $q$ will eventually received $b$, a contradiction.
@@
[Sequential Advance] If round $r$ is advanced in a blocklace $B$, then every round $0<r^{\prime}<r$ is also advanced in $B$.
We next show that the protocol does not get “stuck”, i.e., agents continue to produce blocks when they are needed for progress. Because Grassroots Consensus is quiescent, agents do not have to generate infinitely many blocks. But if some agent issues a block in a round, that round eventually advances.
@@
If a correct agent $p$ sends a block in some round $r$ in Grassroots Consensus, then $r$ eventually becomes advanced at every correct agent.
Proof 3.13.
Let $p_{1}$ be an agent that sends a block $b$ in round $r$. Then when $p_{1}$ does so, $r-1$ is advanced in $p_{1}$’s blocklace.
Assume by way of contradiction that $r$ does not advance at some correct agent $p_{2}$. By Observation 3.11, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace. By Proposition 3.10, this means that no correct agent sends blocks in rounds $>r$.
@@
By Proposition 3.10, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round $r-1$ is advanced in $p_{3}$’s blocklace.
@@
Consider a run of Grassroots Consensus in which the blocklace $B$ of a correct agent $p$ is finite in a suffix of the run. Then $B$ is quiescent.
Proof 3.15.
Let $r$ be the last advanced round in $B$. (Note that round 00 is by definition advanced, so $r$ is always defined). Assume by way of contradiction that $B$ is not quiescent, then $p$ eventually issues a round $r+1$ block $b$. By Proposition 3.12, $r+1$ eventually becomes advanced at $p$. A contradiction.
Proposition 3.16 (Liveness in finite runs).
Consider a run of Grassroots Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all non-empty ordinary blocks sent by correct agents in this run are ordered in $B$.
Proof 3.17.
By Proposition 3.10, $B$ includes all ordinary blocks sent by correct agents. By Proposition 3.14, $B$ is quiescent. This means that its last wave includes a first round block $b$ that is final in $B$, and all other blocks in the wave are empty. Because $b$ is final, $p$ calls $\tau(b)$. Also by assumption of $B$ being quiescent, all non-empty ordinary blocks in $B$ are in $[b]$, and hence ordered by $\tau(b)$.
Next, we show that after GST, all waves that have correct leaders in high throughput mode are finalizing. Because quiescence is not monotonic, we consider here only waves $k$ for which the preceding wave, $k-1$, is not quiescent at any correct agent at any time.
@@
Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round $r-1$ advances at time $t$>GST, and assume that the wave that ends in round $r-1$ is not quiescent in any correct agent’s blocklace at any time. Then $p$ sends a block in round $r$ if and only if it is the round’s leader. Moreover, $p$ endorses a round $r$ leader block in round $r+1$.
Proof 3.19.
Let $p_{\ell}$ be the correct leader of round $r$. If $p=p_{\ell}$ then once $r-1$ is advanced, $p$ immediately issues a leader block for round $r$, causing round $r$ to advance, and issues a block endorsing it in round $r+1$.
Otherwise, $p$ waits for a block from $p_{\ell}$ until time $t+2\Delta$ and then sends $p_{\ell}$ a nudge-block, which is received by $t+3\Delta$, causing $p_{\ell}$ to send $p$ a nack-block and get the blocks it is missing by time $t+5\Delta$. Because round $r-1$ is advanced and its wave is not quiescent, $p_{\ell}$ sends a leader block, which is received by $p$ by time $t+6\Delta$. Some predecessors of this block might be missing, causing $p$ to wait $\Delta$ and then send a nack-block and get the missing blocks allowing it to accept the leader block by time $t+9\Delta$.
@@
Let $r$ be a first round with a correct leader s.t. round $r-1$ advances after GST and the wave that ends in round $r-1$ is not quiescent in any correct agent’s blocklace at any time. Then a round $r$ leader block is finalized in its wave at all correct agents.
Proof 3.21.
@@
Consider a run of Grassroots Consensus in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time. Then all the blocks of correct agents are ordered by all agents in this run.
Proof 3.23.
Let $B$ be the (infinite) union of all blocklaces in the run. Let $b_{1}$ be a block of a correct agent in $B$. By Proposition 3.10, $b_{1}$ is eventually included in all correct agents’ blocklaces. Therefore, it is approved by all correct agents’ blocks from some wave onward. Because every quiescent wave is finalizing, and finalization is monotonic, there are infinitely many finalizing waves in $B$. Let $b_{2}$ be a final block that approves $b_{1}$ in some wave $k$. Once $b_{2}$ is final at some agent’s blocklace, $b_{1}$ is ordered at that agent. We will show that $b_{2}$ is finalized at all agents.
Observe that since $b_{2}$ is finalized in wave $k$, it is approved by a super-majority in wave $k$. Because every valid first round block approves a super-majority in the preceding round and there is at least one correct agent at the intersection of any two super-majorities, every valid first round block in any round $k^{\prime}>k$ observes $b_{2}$.
Consider a finalizing wave $k^{\prime}>k$ in $B$ and let $b_{3}$ be the first-round block that is finalized in wave $k^{\prime}$. Then $b_{3}$ is endorsed by a super-majority in the second round. Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition 3.10). If $b_{3}$ is not a leader block, then to be validly endorsed, $[b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.

### v4

@@ 4.2 Liveness
Next, we discuss liveness. Having proved safety, we know that all output sequences of correct agents are subsequences of a common sequence. Let $T$ be the minimal such sequence. Then the set of elements in $T$ is the union of all the elements in output sequences of correct agents. Because all members of $P$ participate in all epochs, Epoch Liveness requires that all their output sequences include all the transactions in $T$, which are all the transactions output by correct agents. Constitution Amendment Liveness trivially holds in the single-epoch case. We conclude with the following observation:
{observation}
[Single-epoch Liveness] A single-epoch consensus protocol is live if and only if the output sequences of all correct participants include all the transactions input or output by correct participants.
Theorem 4.10.
Single-epoch Constitutional Consensus is live.
We first prove that Constitutional Consensus disseminates all correct agents’ blocks and the blocks they depend on.
@@
If a block $b$ is issued by a correct agent in Constitutional Consensus then $[b]$ is eventually received by all correct agents.
Proof 4.12.
By way of contradiction assume that $p,q\in P$ are correct, $b_{p}$ is issued by $p$ (and received by $q$ since the network is reliable) and some block $b\in[b_{p}]$ is never received by $q$. Since by assumption $b$ (and possibly other blocks in $b_{p}]$) have not been received by $q$, then after receiving $b_{p}$, $q$ issues a nack-block $b_{q}$ for which $b\in[b_{q}]$. Upon receipt of $b_{q}$ by $p$, $p$ will send (once) $b_{q}]$, which includes $b$, to $q$. Since the network is reliable, $q$ will eventually received $b$, a contradiction.
@@
[Sequential Advance] If a blocklace $B$ contains a round $r$ block, then every round $0<r^{\prime}<r$ is advanced in $B$.
We next show that the protocol does not get “stuck”, i.e., agents continue to produce blocks when they are needed for progress. Because Constitutional Consensus is quiescent, agents do not have to generate infinitely many blocks. But if some agent issues a block in a round, that round eventually advances.
@@
If a correct agent $p$ sends a block in some round $r$ in Constitutional Consensus, then $r$ eventually becomes advanced at every correct agent.
Proof 4.14.
Let $p_{1}$ be an agent that sends a block $b$ in round $r$. Then when $p_{1}$ does so, r1r-1 is advanced in $p_{1}$’s blocklace.
Assume by way of contradiction that $r$ does not advance at some correct agent $p_{2}$. By Observation 4.12, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace. By Proposition 4.11, this means that no correct agent sends blocks in rounds >r>r.
@@
By Proposition 4.11, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round r1r-1 is advanced in $p_{3}$’s blocklace.
@@
Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent $p$ is finite in a suffix of the run. Then $B$ is quiescent.
Proof 4.16.
Assume by way of contradiction that $p$’s latest advanced round in the run, $r$, is not quiescent. Then, by the Issue rule, $p$ eventually issues a block in round $r+1$ (at most $9\Delta$ time after $r$ is advanced).
Proposition 4.17 (Block issuance).
If an agent $p$ has payload to send, then $p$ eventually issues a block with this payload.
Proof 4.18.
Assume by contradiction that $p$ has a non-empty payload from some time when $p$’s most advanced round is $r_{0}$ onward, and $p$ never issues a block with this payload. First, consider the case that $p$’s rounds do not advance in some suffix of the run. Then by Proposition 4.15, $p$’s blocklace is quiescent. Hence, condition 1 of the Issue rule is satisfied, and the payload is sent. A contradiction. Therefore, $p$’s advanced rounds increase indefinitely. Consider a time when $p$’s highest advanced round $r>r_{0}$ and the payload has not been sent. Then $p$ issues a block either in round $r+1$ (by the Issue rule) or in round $r$ (by the Issue backlog rule). A contradiction.
Proposition 4.19 (Transaction Liveness in finite runs).
Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all input transactions of correct agents are ordered in $B$.
Proof 4.20.
By Proposition 4.11, $B$ includes all ordinary blocks sent by correct agents. By Proposition 4.15, $B$ is quiescent. This means that its last wave includes a first round block $b$ that is final in $B$, and all other blocks in the wave are empty. Because $b$ is final, $p$ calls $\tau(b)$. Also by assumption of $B$ being quiescent, all non-empty ordinary blocks in $B$ are in $[b]$, and hence ordered by $\tau(b)$. Moreover, by Proposition 4.17, no agents have payloads to send. It follows that all input transactions have been sent in ordinary blocks and ordered.
Proposition 4.21 (Epoch Liveness in finite runs).
Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all transactions output by correct agents are ordered in $B$.
Proof 4.22.
Consider a transaction $t$ output by a correct agent $p_{1}$ and another correct agent $p_{2}$ whose blocklace $B_{2}$ is finite in a suffix of the run. Because $t$ is at some point final in $p_{1}$’s blocklace, its blocklace eventually includes a block $b$ with $t$ in its payload and a super-majority of blocks that observe $b$. This super-majority includes at least one $p_{3}$-block $b^{\prime}$ by a correct agent $p_{3}$. By Proposition 4.11, $p_{2}$ $b^{\prime}$ is included in $B_{2}$, and because $b\in[b]$, so is $b$. Because $B_{2}$ is finite in a suffix of the run, by Proposition 4.15, it is quiescent, which means that all non-empty blocks including $b$ are final in $B_{2}$, as needed.
Next, we show that after GST, all waves that have correct leaders in high throughput mode are finalising. Because quiescence is not monotonic, we consider here only waves $k$ for which the predecessor wave, k1k-1, is not quiescent at any correct agent at any time.
@@
Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round r1r-1 advances at time tt>GST, and assume that the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time. Then $p$ sends a block in round $r$ if and only if it is the round’s leader. Moreover, $p$ endorses a round $r$ leader block in round $r+1$.
Proof 4.24.
Let $p_{\ell}$ be the correct leader of round $r$. If $p=p_{\ell}$ then once r1r-1 is advanced, $p$ immediately issues a leader block for round $r$, causing round $r$ to advance, and issues a block endorsing it in round $r+1$.
Otherwise, $p$ waits for a block from $p_{\ell}$ until time $t+2\Delta$ and then sends $p_{\ell}$ an inform-block, which is received by $t+3\Delta$, causing $p_{\ell}$ to send $p$ a nack-block and get the blocks it is missing by time $t+5\Delta$. Because round r1r-1 is advanced and its wave is not quiescent, $p_{\ell}$ sends a leader block, which is received by $p$ by time $t+6\Delta$. Some predecessors of this block might be missing, causing $p$ to wait $\Delta$ and then send a nack-block and get the missing blocks allowing it to accept the leader block by time $t+9\Delta$.
@@
Let $r$ be a first round with a correct leader s.t. round r1r-1 advances after GST and the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time. Then a round $r$ leader block is finalized in its wave at all correct agents.
Proof 4.26.
@@
Consider a run of Constitutional Consensus in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time. Then all the blocks of correct agents are ordered by all agents in this run.
Proof 4.28.
Let $B$ be the (infinite) union of all blocklaces in the run. Let $b_{1}$ be a block of a correct agent in $B$. By Proposition 4.11, $b_{1}$ is eventually included in all correct agents’ blocklaces. Therefore, it is approved by all correct agents’ blocks from some wave onward. Because every quiescent wave is finalising, and finalization is monotonic, there are infinitely many finalising waves in $B$. Let $b_{2}$ be a final block that approves $b_{1}$ in some wave $k$. Once $b_{2}$ is final at some agent’s blocklace, $b_{1}$ is ordered at that agent. We will show that $b_{2}$ is finalized at all agents.
Observe that since $b_{2}$ is finalized in wave $k$, it is approved by a super-majority in wave $k$. Because every valid first round block approves a super-majority in the previous round and there is at least one correct agent at the intersection of any two super-majorities, every valid first round block in any round $k^{\prime}>k$ observes $b_{2}$.
Consider a finalising wave $k^{\prime}>k$ in $B$ and let $b_{3}$ be the first-round block that is finalized in wave $k^{\prime}$. Then $b_{3}$ is endorsed by a super-majority in the second round. Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition 4.11). If $b_{3}$ is not a leader block, then to be validly endorsed, $b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.